If the current through wire A is 1 A out of the page and the current flowing through wire B is 1 A into the page, what is the magnitude of the magnetic field at point P?

Group of answer choices Im

0.318 x 10-7 T

2 x10-7 T

3.14 x 10-7 T

4 x 10-7 T

1 x 10-7 T
Answer #1

Using right hand thumb rule The magnetic force due to A is upward (+y direction ).

The magnetic force due to B is upward (+y direction).

So resultant field at point P

Bp = BA + BB

Bp = μοIA2πR+ μοΙΒ/2πR

Here No = 41 * 10-7 , IA=IB=1.0 A , R=1.0 M

So

B, = 47 * 10-7 x 1.0/27 * 1.0+ 47 x 10-7 X 1.0/27 * 1.0

B = 2 x 10-7+2 x 10-7 = 4 x 10-7

Bp = 4 x 10-77 Answer